OpenStudy (anonymous):

Please help! Solve for x: 5 − (x + 5) > −2(x + 4) I don't know what to do! ~Micah K.

undeadknight26 (undeadknight26):

would u like to know the answer or how you get the problem?

OpenStudy (anonymous):

Preferably both, but you don't have to give me the answer.

undeadknight26 (undeadknight26):

hmm I know the answer from b4 but i don't remember how i got it...

OpenStudy (anonymous):

That's how it is for me most of the time. I know the answer, but I don't know how to show my work... :(

OpenStudy (anonymous):

try to isolate the x into one side and the numbers into the other side

OpenStudy (anonymous):

On the right side, I think that I have to use the distributive property. Am I right?

OpenStudy (anonymous):

correct

undeadknight26 (undeadknight26):

Lol let me try this... 5 + x = 5x - 5 = 0 x + 4 = 4x * -2 = -8 -8 + 0 = -8 x = -8?

undeadknight26 (undeadknight26):

most likely wrong...but hey! got the right answer right?

OpenStudy (anonymous):

Okay, so that would mean 5 - (x + 5) > -2x - 8

undeadknight26 (undeadknight26):

well no that would mean x > -8

OpenStudy (anonymous):

These are my options: x > −8 x < −8 x > −18 x < −18

undeadknight26 (undeadknight26):

second one!

undeadknight26 (undeadknight26):

no first!

OpenStudy (anonymous):

And yes, you wouldn't use the equal sign. You would use a less or greater than sign.

OpenStudy (anonymous):

the left side would be -x +10

undeadknight26 (undeadknight26):

definitely first one.

OpenStudy (anonymous):

So the answer is A?

OpenStudy (anonymous):

How is the left side -x + 10? @helpme1.2

undeadknight26 (undeadknight26):

I believe so check my awnser again... http://www.wolframalpha.com/input/?i=5+%E2%88%92+%28x+%2B+5%29+%3E+%E2%88%922%28x+%2B+4%29

OpenStudy (anonymous):

you add the 5s which would give u 10

OpenStudy (anonymous):

5s??

OpenStudy (anonymous):

the 2 fives u have 5-(x+5)

OpenStudy (anonymous):

DUDE THANKS FOR THE AWESOME LINK!!!!! :D

OpenStudy (anonymous):

Okay, then what?

OpenStudy (anonymous):

Ohhhhh 5's not 5s.

OpenStudy (anonymous):

I thought that you were saying that "s" was a variable.....

OpenStudy (anonymous):